A positive number is one-half another number. The sum of the numbers is 28 less than the square of the smaller number. What are the numbers?
Answer by ikleyn(53751)

.
A positive number is one-half another number. The sum of the numbers is 28 less
than the square of the smaller number. What are the numbers?
~~~~~~~~~~~~~~~

From the condition, it is clear that BOTH the NUMBERS are positive.

So, let x be the smaller of the two numbers; then the other number is 2x.


Next, write an equation as you read the problem

    x + 2x = x^2 - 28.


It is the same as

     3x   = x^2 - 28,

or

     x^2 - 3x - 28 = 0.


Factor left side

    (x-7)*(x+4) = 0.


The roots are 7 and -4, and because you are looking for positive solutions, 
you accept x= 7 and deny x= -4.


ANSWER.  The numbers are 7 and 14.

You can check it on your own that the answer is correct.

------------

Solved.